Session: Feb 2011 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Wed Mar 02 11:49:24 2011) Minor failures: HIGH
due to heavy snow; Tsys very high on all the Vc.

Torun (Kaz Borkowski at Wed Mar 02 13:39:17 2011) Success: HIGH
Operator noted high Tsys at times. No other known problems.

Effelsberg (Uwe Bach at Wed Mar 02 14:35:21 2011) Minor failures:RECE
Started 10 min too late, because of probelms with frequency settings at the synthesizer.

Yebes (Pablo de Vicente at Wed Mar 02 15:08:42 2011) Success:
No known problems

Westerbork (Geert Kuper at Wed Mar 02 21:24:09 2011) Success:RECE
The last 2 minutes of this project might have bad data due to a receiver steering problem.

Onsala (Michael Lindqvist at Thu Mar 03 07:20:38 2011) Success:
No known problems.

Jodrell_Bank (Alastair Gunn at Tue Mar 29 11:30:15 2011) Success:
Receiver may have been warm. No other known problems.